[Brahms-l] exp update

From: flemming videbaek <videbaek@rcf.rhic.bnl.gov>
Date: Wed Apr 07 2004 - 21:45:41 EDT
As of today, the last day of opening before the run-4 pp running in ernest progress was made allong the following lines

a) The INEL and CC counters were in-coorporate into DAQ and HV; 
b) HV was turned on. So far it looks fine.
b) The T3-T5 survey was completed.
c) T2 DVM measurement were made upstream to checkout if differences were seen.
d) changes has been described in e-log.

Flemming

----------------------------------------------------------------
Flemming Videbaek
Physics Department
Brookhaven National Laboratory

e-mail: videbaek@bnl.gov
phone: 631-344-4106



_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Apr 7 21:39:15 2004

This archive was generated by hypermail 2.1.8 : Wed Apr 07 2004 - 21:39:33 EDT